OpenStudy (anonymous):

PLEASE help me with this! Someone?? /: Write the slope-intercept form of the line parallel to the line in part (a) that passes through (2,0). Graph the line. SHOW ALL WORK! 3. Write the slope-intercept form of the line perpendicular line in part (a) passing through (0, -4). Graph the line. SHOW ALL WORK!

OpenStudy (anonymous):

what is the equation of the line in part (a)??

OpenStudy (anonymous):

y=3x+2

OpenStudy (anonymous):

ok start with the point slope form: \[y-y _{1}=m(x-_{x _{1}}\] now your \[y _{1} and x _{1} \] are going to be the points you were given (0,-4) and the slope of the line perpendicular is the negative reciprocal of the slope in the original equation. Does this help?

OpenStudy (anonymous):

ok so is the answer to 3. y=-1/3x-4 ....??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok thank you!! and is the answer to 2. is y=3(x-2) ?

OpenStudy (anonymous):

what is the question for 2?

OpenStudy (anonymous):

. Write the slope-intercept form of the line parallel to the line in part (a) that passes through (2,0). Graph the line. SHOW ALL WORK!

OpenStudy (anonymous):

yes you are correct, you can also distribute the 3 to get y=3x-6